OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, –2) and has a slope of –3

OpenStudy (anonymous):

is it \[3x+y-10=0\]

OpenStudy (zpupster):

y=mx+b

OpenStudy (zpupster):

(x,y) =( 4,-2)

OpenStudy (anonymous):

-2=-3(4)+b

OpenStudy (zpupster):

usually it is like this: an extended form of the slope intercept equation. y- (-2) = -3(x-4) y+2 = -3x + 12 y= -3x+10

OpenStudy (zpupster):

here is the extended form: y-y1= m(x-x1)

OpenStudy (zpupster):

same as deena's

OpenStudy (anonymous):

3x+y=-10 ?

OpenStudy (zpupster):

you changed the sign on ten

OpenStudy (anonymous):

so = 10 not -10

OpenStudy (zpupster):

yes

OpenStudy (anonymous):

thank you